OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 4 sin2x - 4 sin x + 1 = 0

OpenStudy (ghazi):

just consider sin x= y...you'll have 4 y^2-4 Y+1=0..solve for y first

OpenStudy (anonymous):

y= -4-1

OpenStudy (anonymous):

pi/6 is the first one right?

OpenStudy (ghazi):

nope..solve either by factorizing or by quadratic equation and see what you get \[D= -b \pm \sqrt{b^2-4ac}/2a\]

OpenStudy (ghazi):

yes it's pi/6

OpenStudy (anonymous):

&& then 11pi/6

OpenStudy (ghazi):

nope..then it's 2 pi /5

OpenStudy (ghazi):

just use pi/6 + pi/2 \[\frac{ \pi }{ 6 }+ \frac{ \pi }{ 2 }\]

OpenStudy (ghazi):

@babydoll332 did you get?

OpenStudy (anonymous):

5pi/6

OpenStudy (ghazi):

nope it's 4pi/6= 2pi/3

OpenStudy (anonymous):

i can't find that 1 :/

OpenStudy (ghazi):

take LCM of 2 and 6 you'll have 6 in the denominator and then \[\frac{ \pi }{ 6 }+\frac{ \pi }{ 2 }= \frac{ \pi+3*\pi }{ 6 }= \frac{ \pi +3\pi }{ 6 }= \frac{ 4\pi }{ 6 }= \frac{ 2 \pi }{ 3}\]

OpenStudy (ghazi):

is ti clear?

OpenStudy (ghazi):

see you have pi/6 and you need to find answer between (0,2pi) so just use 2 pi- pi/6